OpenStudy (anonymous):

Lastly, factorize: 5(x-2) - 3(2-x)(x+7)

OpenStudy (anonymous):

hint is that \[2-x=-(x-2)\]

OpenStudy (anonymous):

solution is \[5(x-2)-3(2-x)(x+7)=5(x-2)+3(x-2)(x+7)\] \[=(5+3(x+7))(x-2)\] \[=(5+3x+21)(x-2)\] \[=3x+26)(x+2)\]

OpenStudy (anonymous):

Ty Sat !